A 70N force is applied horizontally to a 10kg block at rest for a displacement of 200m across a frictionless surface. How much work is done by the force? What is the acceleration of the block?


1
Expert's answer
2022-01-10T09:09:07-0500

Given:

"F=70\:\rm N"

"m=10\:\rm kg"

"d=200\:\rm m"


(a) The work done by the force

"W=Fd=70*200=14000\:\rm J"

(b) The acceleration of a body

"a=\frac{F}{m}=\frac{70}{10}=7\:\rm m/s^2"